Future Motors is expected to pay an annual dividend next year of $3.10 a share. Dividends are expected to increase by 1.85 percent annually. What is one share of this stock worth at a required rate of return of 15 percent?


 A) $24.01 B) $26.30 C)$24.56 ) $23.57 ) $24.59
Answer #1

Answer is Option D

This question is based on application of constant growth dividend discount model, according to which

$$ P_{0}=\frac{D i v_{1}}{r-g} $$

\(P_{0}=\) Price of Stock Div \(_{1}=\) Estimated Dividends for Next Period \(r=\) Required Rate of Return \(g=\) Growth Rate

$$ \begin{array}{l} P_{0}=\frac{3.10}{0.15-0.0185} \\ P_{0}=23.57 \end{array} $$